# InterpIoU: Rethinking Bounding Box Regression with Interpolation-Based IoU Optimization

### v1

@@ 1 Introduction
Object detection is a fundamental task in computer vision and plays a vital role in various real-world applications such as intelligent surveillance [23, 7] and autonomous driving [5, 3]. It comprises two primary sub-tasks: object classification and localization. The spatial location of an object is typically represented by a bounding box (Bbox), which encodes its position and shape. The core of localization is Bounding Box Regression (BBR), a training process that aims to maximize the geometric overlap between the predicted Bbox and its ground-truth counterpart. This overlap is universally quantified by the Intersection over Union (IoU) metric. Consequently, the design of an effective localization loss should be fundamentally aligned with the objective of maximizing the IoU value, as this directly reflects localization quality.
@@
Early object detectors such as Fast R-CNN [11] and YOLOv1 [20] employed $\ell_{n}$-norm losses to regress Bbox coordinates. These losses treat the four coordinates independently and ignore the geometric structure of the Bbox. As a result, they are inherently misaligned with the BBR objective and are also sensitive to object scale. To bridge this gap, IoULoss [26] was introduced to directly optimize the IoU metric. However, it suffers from gradient vanishing when the predicted and ground-truth boxes do not overlap, limiting its optimization effectiveness. To address this, a series of extended IoU-based losses—such as GIoU [22], DIoU [31], CIoU [32], SIoU [9], and PIoU [15]—propose augmenting IoU with additional geometric penalty terms (e.g., center distance, aspect ratio, or angle). Despite their empirical improvements, most of these geometric penalties are introduced either to alleviate gradient vanishing in non-overlapping cases or to improve convergence efficiency by providing auxiliary geometric guidance. However, as they are not explicitly designed to maximize IoU, these formulations may impose optimization objectives that diverge from the fundamental goal of BBR. This misalignment can result in unintended behaviors, such as the bounding box enlargement problem [22], where predicted boxes are overly expanded to minimize the auxiliary penalties rather than to genuinely increase overlap with the ground truth. In addition, the effectiveness of these penalty terms often hinges on specific Bbox distributions. Simple geometric terms typically struggle to generalize across diverse object layouts, while more sophisticated ones introduce extra complexity with limited gain. These issues are especially problematic in scenarios involving small or crowded objects [17, 18, 27, 33], where precise and stable gradient signals are essential for accurate localization.
@@ 3.3 Simulation Experiments
In our simulation, we systematically analyzed the influence of distance, scale, and aspect ratio on regression performance. Specifically, we defined seven ground-truth boxes, all having a unit area $S_{gt}$ but with varying aspect ratios: $1\!:\!4,1\!:\!3,1\!:\!2,1\!:\!1,2\!:\!1,3\!:\!1,$ and $4\!:\!1$. The centers of all ground-truth boxes were fixed at $(0.5,0.5)$.
@@ 3.4.1 Visualization of Bbox Enlargement
To illustrate this effect, we designed a simple controlled regression experiment. The center of the initial Bbox was set at $[0.5,0.5]$ with an aspect ratio of $1\!:\!3$, while the target box was located at the origin $(0,0)$ with an aspect ratio of $3\!:\!1$. Both boxes were set to have an area of $1/4$. Typically, the initial anchor box has no intersection with the ground truth box. We used the same optimizer and hyperparameters as in the previous simulation experiments.
@@ 3.4.2 Gradient Analysis
We use $S$ and $S_{g}$ to denote the areas of the prediction Bbox and ground truth Bbox respectively, which can be obtained by $S_{g}=w_{g}h_{g}$ and $S=wh$. We then use $U$ to represent the union area of the two Bboxes: $U=S+S_{g}-I$, where $I$ is the area of the intersection part of the two Bboxes. Next, the basic IoU loss $L{IoU}$ is written as the
@@
For the size parameters $w$ and $h$, the gradients in Eq. (25) and (26) consist of two terms. The first term, $-\frac{S_{g}+S}{U^{2}}\cdot\frac{\partial I}{\partial w}$ (or $-\frac{S_{g}+S}{U^{2}}\cdot\frac{\partial I}{\partial h}$), encourages increasing the intersection area $I$. If increasing $w$ (or $h$) enlarges $I$, the gradient pushes $w$ (or $h$) to grow further. On the other hand, if increasing $w$ (or $h$) reduces $I$, the gradient reverses direction to shrink the box. The second term, $\frac{I}{U^{2}}\cdot\frac{\partial S}{\partial w}$ (or $\frac{I}{U^{2}}\cdot\frac{\partial S}{\partial h}$), penalizes increasing the predicted box area $S$. Since $\frac{\partial S}{\partial w}=h$ and $\frac{\partial S}{\partial h}=w$ are always positive, this term pushes $w$ and $h$ to decrease. Together, these two terms balance the objectives of maximizing $I$ and minimizing $S$.

### v2

@@ 1 Introduction
Object detection is a fundamental task in computer vision and plays a vital role in various real-world applications such as intelligent surveillance [12, 28], industrial automation [8, 6] and autonomous driving [10, 4, 5, 2]. It comprises two primary sub-tasks: object classification and localization. The spatial location of an object is typically represented by a bounding box (Bbox), which encodes its position and shape. The core of localization is Bounding Box Regression (BBR), a training process that aims to maximize the geometric overlap between the predicted Bbox and its ground-truth counterpart. This overlap is universally quantified by the Intersection over Union (IoU) metric. Consequently, the design of an effective localization loss should be fundamentally aligned with the objective of maximizing the IoU value, as this directly reflects localization quality.
@@
Early object detectors such as Fast R-CNN [16] and YOLOv1 [24] employed $\ell_{n}$-norm losses to regress Bbox coordinates. These losses treat the four coordinates independently and ignore the geometric structure of the Bbox. As a result, they are inherently misaligned with the BBR objective and are also sensitive to object scale. To bridge this gap, IoULoss [32] was introduced to directly optimize the IoU metric. However, it suffers from gradient vanishing when the predicted and ground-truth boxes do not overlap, limiting its optimization effectiveness. To address this, a series of extended IoU-based losses—such as GIoU [26], DIoU [37], CIoU [38], SIoU [14], and PIoU [21]—propose augmenting IoU with additional geometric penalty terms (e.g., center distance, aspect ratio, or angle). Despite their empirical improvements, most of these geometric penalties are introduced either to alleviate gradient vanishing in non-overlapping cases or to improve convergence efficiency by providing auxiliary geometric guidance. However, as they are not explicitly designed to maximize IoU, these formulations may impose optimization objectives that diverge from the fundamental goal of BBR. This misalignment can result in unintended behaviors, such as the bounding box enlargement problem [26], where predicted boxes are overly expanded to minimize the auxiliary penalties rather than to genuinely increase overlap with the ground truth. In addition, the effectiveness of these penalty terms often hinges on specific Bbox distributions. Simple geometric terms typically struggle to generalize across diverse object layouts, while more sophisticated ones introduce extra complexity with limited gain. These issues are especially problematic in scenarios involving small or crowded objects [31, 33, 17, 7, 27], where precise and stable gradient signals are essential for accurate localization.
@@ 3.3 Simulation Experiments
In our simulation, we systematically analyzed the influence of distance, scale, and aspect ratio on regression performance. Specifically, we defined seven ground-truth boxes, all having a unit area $S_{gt}$ but with varying aspect ratios: 1:4,1:3,1:2,1:1,2:1,3:1,1\!:\!4,1\!:\!3,1\!:\!2,1\!:\!1,2\!:\!1,3\!:\!1, and 4:14\!:\!1. The centers of all ground-truth boxes were fixed at $(0.5,0.5)$.
@@ 3.4.1 Visualization of Bbox Enlargement
To illustrate this effect, we designed a simple controlled regression experiment. The center of the initial Bbox was set at $[0.5,0.5]$ with an aspect ratio of 1:31\!:\!3, while the target box was located at the origin $(0,0)$ with an aspect ratio of 3:13\!:\!1. Both boxes were set to have an area of $1/4$. Typically, the initial anchor box has no intersection with the ground truth box. We used the same optimizer and hyperparameters as in the previous simulation experiments.
@@ 3.4.2 Gradient Analysis
We use $S$ and $S_{g}$ to denote the areas of the prediction Bbox and ground truth Bbox respectively, which can be obtained by $S_{g}=w_{g}h_{g}$ and $S=wh$. We then use $U$ to represent the union area of the two Bboxes: $U=S+S_{g}-I$, where $I$ is the area of the intersection part of the two Bboxes. Next, the basic IoU loss LIoUL{IoU} is written as the
@@
For the size parameters $w$ and $h$, the gradients in Eq. (25) and (26) consist of two terms. The first term, $\frac{S_{g}+S}{U^{2}}\cdot\frac{\partial I}{\partial w}$ (or $\frac{S_{g}+S}{U^{2}}\cdot\frac{\partial I}{\partial h}$), encourages increasing the intersection area $I$. If increasing $w$ (or $h$) enlarges $I$, the gradient pushes $w$ (or $h$) to grow further. On the other hand, if increasing $w$ (or $h$) reduces $I$, the gradient reverses direction to shrink the box. The second term, $\frac{I}{U^{2}}\cdot\frac{\partial S}{\partial w}$ (or $\frac{I}{U^{2}}\cdot\frac{\partial S}{\partial h}$), penalizes increasing the predicted box area $S$. Since $\frac{\partial S}{\partial w}=h$ and $\frac{\partial S}{\partial h}=w$ are always positive, this term pushes $w$ and $h$ to decrease. Together, these two terms balance the objectives of maximizing $I$ and minimizing $S$.